# Euclid Quick Data Release (Q1)

### v1

@@
Ever since the discovery that the mass or dynamics of a spheroid within a galaxy correlates with the mass of any detectable supermassive black hole (SMBH) at its centre (e.g., Dressler 1989; Kormendy & Richstone 1995; Magorrian et al. 1998; Ferrarese & Merritt 2000, and more recently, Sahu et al. 2019; Davis et al. 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre. If galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation leading it to be classified as an active galactic nucleus (AGN). Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
@@
For imaging in wavebands such as the optical and near-IR, the presence and detectability of emission from the area directly around a SMBH depends on multiple factors. This can include the rate of infall of material into the region around the SMBH, the amount of obscuration (absorption or scattering) towards that region, the contrast between the SMBH-related emission and that of the surrounding galaxy, the redshift, and the intrinsic compactness of the stellar distribution of the host galaxy. The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review). Consequently, identifying and characterising SMBHs, particularly during periods of active feedback is central to the study of SMBH and galaxy evolution.
@@
A more precise classification of a source as an AGN is possible by using a much wider range of multi-wavelength data. This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN. However, it is still worth considering the efficient selection of AGN candidates from single-band imaging as a first step in identifying large samples of AGN for further study. In particular, with the advent of the uniform, deep and high spatial-resolution optical imaging data set from the Euclid mission, covering a large fraction of the extragalactic sky, there is the potential to derive unprecedentedly large samples of AGN across a wide range of (optical) luminosity and redshift parameter space with the use of an appropriate and sufficiently efficient technique. The technique can then be subsequently refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates. With sufficient refinement, this further assessment step subsequently becomes either unnecessary or at the very least an efficient stage in compiling reliable samples. This then opens the door to production of unprecedentedly large and reliable samples of optically-selected AGN.
@@
In this work, we present a technique that can achieve the above goal. In the following we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS band (Euclid Collaboration: Mellier et al. 2024; Euclid Collaboration: Cropper et al. 2024) imaging alone. Diffusion-based methods are machine-learning models used to generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps. This technique outperforms traditional methods like median pixel substitution or interpolation, which often struggle to recreate complex structures within images. By using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels. Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
Briefly, the technique takes an image of a galaxy that may or may not have AGN emission at its centre and excludes those central pixels that may be influenced by AGN emission. The model then predicts (‘inpaints’) the underlying stellar light profile in those pixels from a comparison of the rest of the galaxy image to a model derived from a very large sample of similar galaxy images. Any significant difference between the prediction and true image in those central regions indicates the potential presence of AGN emission. Because AGN are comparatively rare, the sample of galaxy images used in the generation of the model does not need to be filtered to reject AGN, the method will be inherently biased to the ‘typical’ behaviour of galaxy light distributions and so naturally down-weights or excludes predicting central regions with AGN emission.
@@
While this paper details the basic technique applied to VIS imaging data, as this is a machine-learning technique, future enhancements would allow incorporation of the results of a detailed assessment step, involving other data sets, in order to further improve the accuracy and reliability of the AGN candidate samples produced directly by the technique. Such steps could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps. Additionally, the technique as presented here uses a standard treatment of diffusion models which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data. Such work would be the subject of future papers.
@@ 2.1 Generative models
Generative models such as generative adversarial networks (GANs, Goodfellow et al. 2014, 2020) or variational autoencoders (VAEs, Kingma 2013) are trained with the aim to learn the underlying distribution of a data set to generate novel examples.
Using a pair of competing networks, GANs ensure that once trained the generated images are of high quality, leading to the creation of many artificially generated data sets that are indistinguishable from the original data set (Karras et al. 2020). However, the adversarial training loop is often difficult to train, often leading to mode collapse, meaning the generator focuses on only a small subset of the total search space leading to generating samples of similar instances (Thanh-Tung & Tran 2020).
On the other hand, VAEs use an encoder network to compress the data into a latent space of fewer dimensions. The network then uses the latent space to recreate the input through an often symmetric decoder network. VAEs provide better coverage over the search space, improving stability during training as well as increased variety in generated outputs. However, the drawback for VAEs becomes apparent when handling more complex, high dimensional data. The use of latent spaces does not guarantee a lossless compression of information, often leading to accurate but blurry outputs.
@@ 2.2 Diffusion models
Diffusion-based models provide an alternative process for image generation that allows for both good coverage, as well as the ability to represent complex data distributions (Dhariwal & Nichol 2021). For both GANs and VAEs, their significant advantage with respect to diffusion models is that they offer faster sample generation and lower computational costs. Whereas diffusion models trade off speed for superior sample quality, mode coverage, and flexibility in controllable generation making them our preferred choice in the task of AGN identification.
This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al. 2020). Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually one is left with an image of complete noise without any remnants of the original input. If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process. By predicting the noise that was added to an image at timestep , we can remove this noise to produce image . This forms an iterative process originating from some final timestep , an image of pure noise, to a realistic image from the data distribution at timestep .
@@
The term defines the variance of the Gaussian noise added at each timestep. By ensuring a gradual increase in noise, the model is more easily able to learn the transition from the data distribution to the noise distribution.
@@
Smith et al. (2022) also utilised the DDPM framework to generate a data set of realistic galaxies. The model is trained using the Photometry and Rotation curve OBservations from Extragalactic Surveys (PROBES) data set (Stone & Courteau 2019; Stone et al. 2021), a collection of large, well resolved objects that feature significant internal structure. They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create not only visually realistic images, but also how they feature similar physical property distributions such as half-light radius and flux-space colour values. They also briefly explore the use of inpainting with their model to remove satellite trails.
@@
When researching diffusion models, papers utilising score-based models are often used and spoken of interchangeably, as they are different implementations of the same generative process (Smith et al. 2022). Rather than the fixed sequence of timesteps to denoise data, score-based methods apply stochastic differential equations (SDEs) to estimate the data distribution’s gradient (score). Working with SDEs and gradients allows for a continuous range of possible diffusion paths. It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models. The use of score-based generative models within astronomical applications include galaxy image deconvolution (Adam et al. 2023; Spagnoletti et al. 2024), gravitation lensing analysis (Adam et al. 2022; Remy et al. 2023), and deblending (Sampson et al. 2024).
@@ 2.3 Noise scheduling
The larger the number of total timesteps , the more refined the model can become at accurately generating realistic images, since less complex noise dynamics are required to be learnt for each timestep . This does, however, require the model to learn the dynamics of more timesteps, leading to extra overhead in both training and inference, since more steps must be performed to traverse from pure noise to realistic images.
@@
In the original DDPM paper, Ho et al. (2020) made use of a linear schedule for the amount of noise added to the image at each . Although the produced images were competing with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low resolution images and proposed using a cosine-beta noise schedule as a replacement. By stretching the noise levels so that less noise is added in early iterations and more in the later, the process overcame the issue of images became too noisy too quickly. Figure 2 shows how the scheduler impacts the parameter used in Eq. 2. Equation 3 shows how is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation:

### v2

@@
Ever since the discovery that the mass or dynamics of a spheroid within a galaxy correlates with the mass of any detectable supermassive black hole (SMBH) at its centre (e.g., Dressler 1989; Kormendy & Richstone 1995; Magorrian et al. 1998; Ferrarese & Merritt 2000, and more recently, Sahu et al. 2019; Davis et al. 2018, 2019 and references therein), it has been believed that most – if not all – galaxies with a spheroidal (or pseudo-spheroidal) component contain a SMBH at their centre. If a galaxy’s emission shows evidence for its central SMBH being actively fuelled, it will emit non-thermal radiation, leading it to be classified as an active galactic nucleus (AGN). Depending upon the luminosity of that emission and its contrast with the stellar emission from the surrounding galaxy, it could have a classification ranging from a Seyfert galaxy (for the weakest emission) through to a quasi-stellar object or QSO (for the most luminous).
@@
For imaging in wavebands such as the optical and near-IR, the presence and detectability of emission from the area directly around a SMBH depends on multiple factors. This can include the rate of infall of material into the region around the SMBH, the amount of obscuration (absorption or scattering) towards that region, the contrast between the SMBH-related emission and that of the surrounding galaxy, the redshift, and the intrinsic compactness of the stellar distribution of the host galaxy. The growth and mass of a central SMBH is intimately related to the evolution of its host galaxy, not least through the action of energy, momentum and radiative feedback from outflows originating from the area around the SMBH on the properties of the stellar populations and ISM (see Harrison & Ramos Almeida 2024 for a recent review). Consequently, identifying and characterising SMBHs, particularly during periods of active feedback, is central to the study of SMBH and galaxy evolution.
@@
A more precise classification of a source as an AGN is possible when a much wider range of multi-wavelength data is available. This can include spectroscopy and considering the much broader spectral energy distribution of the galaxy and/or the candidate AGN. However, it is still worth considering the efficient selection of AGN candidates from single-band imaging as a first step in identifying large samples of AGN for further study. In particular, with the advent of the uniform, deep and high spatial-resolution optical imaging data set from the Euclid mission, covering a large fraction of the extragalactic sky, there is the potential to derive unprecedentedly large samples of AGN across a wide range of (optical) luminosity and redshift parameter space with the use of an appropriate and sufficiently efficient technique. The technique can then be subsequently refined by feeding back the result of using other multi-wavelength data, allowing us to confirm or reject candidates. With sufficient refinement, this further assessment step eventually becomes either unnecessary, or at the very least, an efficient stage in compiling reliable samples. This then opens the door to the production of unprecedentedly large and reliable samples of optically-selected AGN.
@@
This work presents a technique that can achieve the above goal. In the following, we describe the use of diffusion models for generating samples of AGN candidates using Euclid VIS band (Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Cropper et al. 2025) imaging alone. Diffusion-based methods are machine-learning models that generate or reconstruct data by evolving a random distribution of pixels into a structured output over several steps. This technique outperforms traditional methods like median pixel substitution or interpolation, which often struggle to recreate complex structures within images. Using these models for ‘inpainting’, where we adaptively recreate parts of the image, we can measure the discrepancies between the original and generated pixels. Examining these errors allows us to separate AGN, QSOs, and other sources from the broader galaxy population.
Our method identifies potential active galactic nuclei (AGN) by treating the task as an unsupervised anomaly detection problem. It begins with an image of a galaxy, potentially containing AGN emission at its center, and masks out the central pixels that might be influenced by such emission. A diffusion model, trained on a large, unlabeled data set of galaxies – most of which are normal, non-active – then predicts or ‘inpaints’ the masked region based on the learned prior distribution of typical galaxy morphologies. Because AGN are rare, there is no requirement to remove them from the training sample; the model is inherently biased toward reconstructing the central region as if it belonged to a typical, non-active galaxy. As a result, any significant discrepancy between the inpainted and actual central region suggests the presence of an AGN. This reconstruction error thus serves as a powerful anomaly score for identifying candidate AGN.
@@
While this paper details the fundamental technique applied to VIS imaging data, as this approach utilises machine learning, future improvements could incorporate findings from a detailed assessment that involves additional data sets. These could include multi-wavelength cross-matching or spectroscopic validation, from which we can ultimately reduce the need for further time-expensive refinement steps. Such enhancements would help to further improve the accuracy and reliability of the AGN candidate samples generated by the technique. Additionally, the technique as presented here uses a standard treatment of diffusion models, which in principle can be significantly further optimised for the specific characteristics of the Euclid data set, thereby improving further the efficiency of the processing of the Euclid data. Such work would be the subject of future papers.
@@ 2.1 Generative models
Generative models are a class of machine learning systems designed to learn an underlying data distribution, , from a finite set of independent and identically distributed (i.i.d.) samples . The primary goal of these models is to develop a tractable method for generating new, synthetic samples that are statistically indistinguishable from the training data. Over the years, several distinct families of generative models have emerged, each with unique architectural and training paradigms. A key distinction lies in their optimisation objectives, which often seek to minimise a measure of divergence between the true data distribution and the model’s learned distribution . The choice of divergence has significant implications for model behavior. Prominent examples include Variational Autoencoders (VAEs, Kingma 2013), which learn a latent variable model of the data, and Generative Adversarial Networks (GANs, Goodfellow et al. 2014, 2020), which involve a contest between a generator and a discriminator. GANs implicitly optimise an objective related to the reverse Kullback-Leibler (KL) divergence, , which exhibits a ‘mode-seeking’ property. This encourages the generation of high-fidelity samples but can lead to mode collapse, where the model learns only a subset of the data distribution (Thanh-Tung & Tran 2020). In contrast, VAEs and Diffusion Models optimise objectives related to the forward KL divergence, , which is ‘mass-covering’. This property encourages the model to account for all modes present in the data, leading to greater sample diversity and more comprehensive coverage of the data distribution.
@@ 2.2 Diffusion models
Diffusion Models (Sohl-Dickstein et al. 2015; Ho et al. 2020), the focus of this work, systematically corrupt data with noise and then learn to reverse the process. They have been shown to achieve superior sample quality and mode coverage compared to other classes of models (Dhariwal & Nichol 2021), making them a compelling choice for modelling the complex and diverse morphologies of galaxies observed by Euclid.
For both GANs and VAEs, their significant advantage compared to diffusion models is that they offer faster sample generation and lower computational costs. Conversely, diffusion models trade off speed for superior sample quality, mode coverage, and flexibility in controllable generation, making them our preferred choice in the task of AGN identification.
This paper makes use of one such implementation, denoising diffusion probabilistic models (DDPMs, Ho et al. 2020). Diffusion works on the following premise: by continually adding a small amount of noise to an image, eventually, one is left with an image of complete noise without any remnants of the original input. If the noise is generated in a stochastic but consistent way, such as from a Gaussian distribution, we can use a network to learn the dynamics of the noise and perform a reverse process. By predicting the noise that was added to an image at timestep , we can remove this noise to produce image . This forms an iterative process originating from some final timestep , an image of pure noise, to a realistic image from the data distribution at timestep .
@@
The term defines the variance of the Gaussian noise added at each timestep. By ensuring a gradual increase in noise, the model can more easily learn the transition from the data distribution to the noise distribution.
@@
Background noise versus diffusion noise
It is important to clarify the distinctions between the definition and use of ‘noise’ in diffusion models compared to the traditional understanding of noise in astronomy-based data. In the context of astronomical imaging, noise refers to the random fluctuations inherent in observational data, often caused by e.g. background emission, electronics read-out noise, and atmospheric disturbances. This ‘background noise’ reduces the signal-to-noise ratio (S/N) of the image, affecting the ability to detect subtle details in any analysis.
In contrast, the diffusion process uses an intentional integration of Gaussian noise to facilitate the training of the generative model. The pixels of image , follow the distribution: , rather than following the distribution of the background noise. The goal of the reverse diffusion process is to remove the artificial, high-variance Gaussian noise that was intentionally added during the forward process, thereby transforming a sample from a simple prior distribution () into a realistic galaxy image. A realistic galaxy image, as defined by our training data set, inherently includes the source’s light profile superimposed on the characteristic background noise of the Euclid instrument. Therefore, the model learns to generate images that replicate this entire distribution, including the statistical properties of the background noise. The diffusion pipeline does not aim to produce a ‘denoised’ image in the traditional astronomical sense (i.e., a background-subtracted image).
Use in Astronomy Applications
@@
Smith et al. (2022) also utilised the DDPM framework to generate a data set of realistic galaxies. The model is trained using the Photometry and Rotation curve OBservations from Extragalactic Surveys (PROBES) data set (Stone & Courteau 2019; Stone et al. 2021), a collection of large, well resolved objects that feature significant internal structure. They provide analysis of the quality of their generated images, showcasing the diffusion model’s ability to create visually realistic images which feature similar physical property distributions such as half-light radius and flux-space colour values. They also briefly explore the use of inpainting with their model to remove satellite trails.
@@
When researching diffusion models, papers utilising score-based models are often used and spoken of interchangeably, as they are different implementations of the same generative process (Smith et al. 2022). Rather than the fixed sequence of timesteps to denoise data, score-based methods apply stochastic differential equations (SDEs) to estimate the data distribution’s gradient (score). Working with SDEs and gradients allows for a continuous range of possible diffusion paths. It is for this reason that score-based models can be classed as the more general framework for diffusion-based generative models. The use of score-based generative models within astronomical applications includes galaxy image deconvolution (Adam et al. 2023; Spagnoletti et al. 2024), gravitation lensing analysis (Adam et al. 2022; Remy et al. 2023), and deblending (Sampson et al. 2024).
2.3 Training Objective
For training, we use 6464 pixel VIS cutouts for each of the selected sources that will be discussed in Sect. 4.1. We use the training pipeline described in Dhariwal & Nichol (2021)111Available at https://github.com/openai/guided-diffusion and adopt a hybrid loss function that combines a simple mean squared error (MSE) term with the full variational lower bound (VLB).
The MSE objective, proposed by Ho et al. (2020), is an MSE loss between the true and predicted Gaussian noise,
where is an original image, is a uniformly sampled timestep, and is the true noise. For our implementation is described below, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
While this objective is effective for training the mean of the reverse process transition, , it provides no learning signal for its variance, . To address this, Nichol & Dhariwal (2021) proposed learning this variance by incorporating the full VLB of the log-likelihood into the training objective. The VLB is given by
Since explicitly depends on the parameters of , it can be used to train them. We adopt the hybrid objective
We use a small weighting constant () to prevent the term from overpowering the term, which remains the primary driver of sample quality. This hybrid loss allows the model to learn an optimal reverse process variance, which can improve model likelihoods and enable more efficient sampling. A key challenge with astronomical data is its high dynamic range. A standard MSE-based loss is not scale-invariant and will be dominated by the few brightest pixels or sources in a training batch. This can lead to poor performance on the much larger population of fainter objects. To address this, we introduce a normalised version of the MSE loss term. For each image , we divide its error by its respective maximum pixel value, denoted as . This normalisation ensures that the relative error is penalised, allowing for more equitable training across both bright and faint sources. The final loss used for our primary model is thus
Throughout the remainder of the paper, we refer to this variant of the loss as the normalised loss. The training curves of this scale-invariant loss compared with the original loss used in Nichol & Dhariwal (2021) is shown in Fig. 18. By incorporating this normalised loss, we enable more equitable training across the full range of galaxy brightnesses, rather than being dominated by the brightest objects which occurs when using the default loss.
@@ 2.4 Noise scheduling
The total number of timesteps, , is a critical hyperparameter in the diffusion process. A larger corresponds to smaller, more refined steps in both the forward (noising) and reverse (denoising) processes. While this can, in principle, allow the model to learn a more accurate representation of the data distribution, it comes with a significant trade-off. The computational cost of inference (i.e., generating a sample) is directly proportional to , as the reverse process must be iterated times. The cost of a single training step, however, is not directly dependent on , because at each step a random timestep is sampled and the loss is evaluated only for that level of noise.
@@
In the original DDPM paper, Ho et al. (2020) made use of a linear schedule for the amount of noise added to the image at each . Although the produced images competed with state-of-the-art methods, Nichol & Dhariwal (2021) found that a linear schedule was not well suited to low-resolution images and proposed using a cosine-beta noise schedule as a replacement. By stretching the noise levels to add less noise in early iterations and more in the latter, the process overcame the issue of images becoming too noisy too quickly. Figure 2 shows how the scheduler impacts the parameter used in Eq. 2. Equation 10 shows how is calculated, where the reader is directed to Nichol & Dhariwal (2021) for specifics on its derivation
where is shown in Eq. 4.
All models in this paper make use of the cosine-beta scheduler. The impact of this scheduler on the high dynamic range nature of astronomy data is explored in detail in Sect. 6.1.
2.5 Conditioning the model
Using the pipeline of DDPMs allows us to impose conditioning on our input space. Throughout training, the model is already conditioned on the timestep . This conditioning is crucial, as it informs the neural network of the noise level present in the input image . The network learns a single function, , that is trained to predict the noise component from a noised image at any given timestep . The characteristics of the noise to be removed differ substantially between a timestep close to (where the image is almost pure noise) and a timestep close to (where the image is almost fully reconstructed). Providing as an input allows the single network to adapt its predictions accordingly.
However, we are not limited to only conditioning by timestep; often, multiple parameters are provided alongside each image. Typically, this is to allow ‘class-conditioning’, where users can tune the expected output of the diffusion model to a specific class or category (Radford et al. 2021; Zhang et al. 2023). By providing an embedding space during training, the model can learn not only the dynamics of the noise but also the inherent differences between noise and non-noise pixel values for each of the class representations. This conditioning provides small nudges to the random walk, guiding the denoising to the respective population of the class. Providing rich representations to the model has led to the impressive text-to-image capabilities of popular diffusion-based implementations, such as Dall-E (Ramesh et al. 2022) or Imagen (Saharia et al. 2022).
This type of conditioning allows the user to direct the image generation to a particular class or style of image. However, with diffusion’s ability to cover large areas of the search space, a repeat sample of the same prompt will likely produce an image very different from the first, even if the generated image belongs to the requested class. Even with the possibility of detailed prompts and refined classes, which can guide the generation towards a narrower or more specific region of the learned space, it may be difficult to know exactly what the output image will look like. This is due to the one-to-many relationship between classes/prompts and their subsequent generated outputs.
However, for our task, we require a different form of conditioning. Instead of generating an entire image based on a class label, we need to reconstruct a specific, masked region of an existing image, conditioned on the surrounding, unmasked pixels. This task is known as inpainting (Pathak et al. 2016; Yeh et al. 2017; Yu et al. 2018; Wang et al. 2018). To achieve this, we do not train a new, specialised conditional model. Instead, we employ a ‘plug-and-play’ (PnP) approach, where conditioning is introduced during the inference phase of an unconditionally trained generative model (Nguyen et al. 2017; Graikos et al. 2022). Our diffusion model, trained on a vast sample of galaxies, serves as a powerful, non-parametric prior for galaxy morphology. The PnP approach allows us to apply this prior to the inpainting task with great flexibility, as the model does not need to be retrained for different images or mask geometries. This is particularly advantageous for large-scale surveys like Euclid. Recent works in astrophysics have successfully employed similar PnP or posterior sampling strategies for complex inference tasks (Feng et al. 2023; Dia et al. 2025).
2.6 Repainting
For this work, we use the Repaint algorithm (Lugmayr et al. 2022), a PnP method specifically designed for diffusion models. The core challenge in diffusion-based inpainting is to ensure semantic and structural consistency between the known (unmasked) and unknown (masked) regions. A naive approach of simply replacing the masked region with noise and then running the reverse process would fail, as the denoising network would alter the known, unmasked pixels, corrupting the conditioning information.
Repaint overcomes this by iteratively enforcing the data constraint at each step of the reverse diffusion process. Given an image and a binary mask (where indicates the region to be inpainted and indicates the known region), the algorithm proceeds as follows. At each timestep of the reverse process, starting from a noisy sample :
1.
Sample the known region: The original image, , is noised to the appropriate level for timestep using the forward process , resulting in a sample of the known part, .
2.
Sample the unknown region: The output image from the previous, slightly more noisy step, , is denoised for one step using the unconditional diffusion model to get a sample of the unknown part, .
3.
Combine and denoise: The known and unknown parts are combined: . This composite image is then passed to the next iteration of the repaint loop to produce the estimate for step . Once this image has been passed through one step of the learned reverse diffusion process , it will produce the slightly less noisy image .
This procedure, illustrated in Fig. 1, ensures that the generated region is coherently conditioned on the surrounding galaxy structure at every stage of the generation. The ‘denoise’ operation in this context refers to the application of one step of the learned reverse diffusion process, , which involves using the trained neural network to predict the noise in and deriving an estimate for . By denoising an image composed of both fixed and generated pixels – each matched to the appropriate noise level – we ensure that the generated region is properly conditioned on the information from the rest of the galaxy, resulting in a final image where all pixels are consistent and coherent.
While alternative conditioning methods based on guiding the reverse process with a likelihood score function exist (Graikos et al. 2022), we adopt the Repaint algorithm for its straightforward implementation and proven effectiveness in enforcing data consistency in a plug-and-play setting.
Incorporating the Repaint method for AGN identification has the following benefits:
Diffusion methods have proven their ability in many applications to accurately recreate complex data sets, allowing for the generation of realistic galaxy morphology images.
A large imbalance of normal galaxies exists compared to AGN, allowing the model to learn a bias. Such a bias can be exploited to create an outlier detection-based classifier without explicitly requiring labels during training.
There are no constraints on the size and shape of masks used, and since Repainting is only applied at inference, no additional retraining is required.
3 Training and inference pipeline
The diffusion and inpainting methods are two distinct pipelines, one for the initial training of the model and a second for the actual inpainting. Training the diffusion model uses the whole image and learns to predict what noise to remove at each timestep, allowing for generating novel examples from the trained distribution.
3.1 Mask creation
Due to the nature of the application, our aim is to mask out any suspected AGN without the need to rely on labels ahead of time. Each mask is positioned so that the brightest pixel in the centre of the source becomes the centre of the mask. Although each image is centred on the source, various interactions and dynamics allow suspected AGN to be offset from the centre. To allow for some deviation, a 99 pixel window around the centre of the image is used to detect the brightest pixel. This offset allows for the correct masking of less symmetric sources whilst limiting the likelihood of masking an adjacent source by mistake. The masks for each image are assigned automatically throughout the data-loading pipeline.
Throughout this paper, we use a 55 pixel mask, unless explicitly stated otherwise, centred on the calculated brightest pixel. A 55 pixel mask, which covers 0505, allows us to effectively cover the AGN contribution whilst minimising loss of the surrounding galaxy structure. Given that the VIS PSF has a full width half maximum (FWHM) of 013 (Euclid Collaboration: Mellier et al. 2025), the AGN’s light is spread over multiple pixels. Taking into account the redshift of the galaxy, AGN within nearby sources () will cover 4.8 pixels, ensuring that we can fully encapsulate the core without unnecessary overlap of the host morphology. By retaining as much of the host galaxy information as possible for inpainting, we also reduce the risk of introducing artificial features into the reconstructed image.
3.2 Reconstruction rescaling
Unlike some image processing pipelines that rescale image data to a limited range, such as when working with JPG images which are binned to the standard 0–255 pixel scale range and converted to floats between 0–1 (Walmsley et al. 2020; Euclid Collaboration: Aussel et al. 2024), our approach utilises the raw values direct from the sensor as input to our network. As comparisons of individual pixels are being made on such a small mask, the relative differences become much more impactful and would likely be lost if values were binned.
One potential issue with this setup is that the libraries used for training are optimised for images within the standard pixel range. Although this did not prevent the model from effectively learning the distribution, it does require a rescaling of the generated output as the pixels are mostly between 1 and 1 to improve training stability.
We experimented with simply normalising the input data so that all our images were in the range 0–1; however, due to the large dynamic range of the data, the faintest pixels ended up being . Given the previous discussions on how MSE does not penalise smaller errors effectively, the model could not accurately recreate sources, outputting a near-empty image of very low pixel values.
To allow the inpainted region to be meaningfully compared (e.g., via MSE) to the original pixel values in that masked region, a post-inference rescaling must be applied. The vast majority of pixels in the image (the unmasked region, which constitutes 99.5% of a 6464 image when a 55 central mask is used) correspond to known pixel values from the original astronomical image. We leverage these known, unmasked pixels to perform a linear rescaling of the diffusion model’s initial output (which is in a normalised range) back to the original image’s dynamic range. This procedure ensures consistency in the unmasked regions and appropriately scales the newly inpainted pixels in the masked region. The following scaling can be applied instantly on output of the repaint procedure using each image’s respective values to allow the non-masked pixels of the input and output images to be consistent: